# Waste Heat and Habitability: Constraints from Technological Energy Consumption

### v1

@@ Abstract
Waste heat production represents an inevitable consequence of energy conversion as per the laws of thermodynamics. Based on this fact, by using simple theoretical models, we analyze constraints on the habitability of Earth-like terrestrial planets hosting putative technological species and technospheres characterized by persistent exponential growth of energy consumption and waste heat generation: in particular, we quantify the deleterious effects of rising surface temperature on biospheric processes and the eventual loss of liquid water. Irrespective of whether these sources of energy are ultimately stellar or planetary (e.g., nuclear, fossil fuels) in nature, we demonstrate that the loss of habitable conditions on such terrestrial planets may be expected to occur on timescales of $\lesssim 1000$ years, as measured from the start of the exponential phase, provided that the annual growth rate of energy consumption is of order $1\%$. We conclude by discussing the types of evolutionary trajectories that might be feasible for industrialized technological species, and sketch the ensuing implications for technosignature searches.
@@ 1 Introduction
It is now firmly established that the upsurge in atmospheric CO2 concentration caused by human technological activities has been elevating Earth’s average surface temperature (e.g., Archer, 2016; Hartmann, 2016; Dessler, 2021; Magnan et al., 2021; Abbass et al., 2022; Romm, 2022; Tziperman, 2022). Nevertheless, it is worth appreciating that ongoing anthropogenic warming is not exclusively a consequence of the intensified greenhouse effect stemming from fossil fuel utilization (though the latter phenomenon is unequivocally proven to be the chief driver). Instead, as elucidated in a number of analyses and overviews (e.g., Chaisson, 2008; Flanner, 2009; Kleidon, 2012; Ming et al., 2014; Lu et al., 2017; Kleidon et al., 2023), this facet embodies a broader outcome of energy production and conversion underpinning modern (viz., globalized and industrialized) human societies. The pioneering 1969 publication by the Soviet scientist Mikhail I. Budyko explicitly underscored this theme (Budyko, 1969, pg. 618):
@@
All the energy used by man is transformed into heat, the main portion of this energy being an additional source of heat as compared to the present radiation gain. Simple calculations show (Budyko 1961) that with the present rate of growth of using energy the heat produced by man in less than two hundred years will be comparable with the energy coming from the sun.
@@
This feature arises from the fundamental fact that energy transduction inevitably generates waste heat (e.g., Budyko, 1969; Summers, 1971; Smil, 2008) [^1] Interestingly, around the same time as Budyko’s paper, Larry Niven’s classic science fiction novel Ringworld (1970) also acknowledged the importance of waste heat as an evolutionary factor for technological species (Niven, 1970).. Although the current contribution of this factor to global warming on Earth is negligible compared to the impact of carbon emissions (Ming et al., 2014; Kleidon et al., 2023), [^2] In localized regions endowed with high industrialization (e.g., large and dense cities), however, it is already a notable contributor to warming (Ichinose et al., 1999; Block et al., 2004; Ohashi et al., 2007; Allen et al., 2011; Zhang et al., 2013; Bohnenstengel et al., 2014; Salamanca et al., 2014; Stewart and Kennedy, 2017; Sun et al., 2018; Raj et al., 2020; Molnár et al., 2020), and this effect is anticipated to increase in the future. it is expected to amplify quite substantially in the next century in the event that the current growth rate of global energy consumption persists (Lu et al., 2017); the consequent rise in global temperature (just from this process) derived from assorted climate models may reach up to $\sim 1$ K in this duration (Flanner, 2009; Nickolaenko, 2009; Steiglechner et al., 2019).
@@
Cruder estimates, when extrapolated over longer intervals, project a temperature increase of $\sim 10$ K over a period of $350$ years, upon assuming similar energy trends as today (Chaisson, 2008). Likewise, a cognate analysis by Mullan and Haqq-Misra (2019) (see also von Hoerner 1975; Dutil and Dumas 2009) concluded that direct heating engendered by progressively higher energy consumption consistent with current rates could dramatically impede human habitability on timescales of merely hundreds of years in the future. It is crucial to realize that this contribution to global warming would persist even in the scenario of achieving a complete transition to zero CO2-emissions energy production.
@@
While this effect is patently of profound importance for the long-term future and fate of human civilization(s), it also has broader significance in the general context of planetary evolution and the potential for harboring technologically capable life in the universe – this latter domain is closely linked with detecting technosignatures, a burgeoning area of astrobiology (Ćirković, 2012; Lingam and Loeb, 2021; Socas-Navarro et al., 2021), with roots traceable to the 1960s (Sagan, 1963; Drake, 1965; Shklovskii and Sagan, 1966; Tarter, 2001). To be specific, we are free to inquire whether the energetic growth, often innately presumed to be associated with technological activity (Kardashev, 1985), imposes a universal fundamental limit on the habitability of a planetary environment, arising simply from thermodynamic considerations, as indicated in von Hoerner (1975).
@@
Thus, the summum bonum of this paper is to perform a preliminary investigation of how the habitability of Earth-like planets would be affected by the presence of life specifically capable of engaging in progressively energy-demanding technological activities. One of our primary objectives in this vein is to gauge the maximum lifespan of habitable planets that harbor an evolving “technosphere” (Donges et al., 2017; Zalasiewicz et al., 2017; Herrmann-Pillath, 2018), by making certain assumptions regarding the growth of energy consumption by an industrialized technological species. This analysis parallels the canonical approach used to estimate the lifetime of continuously habitable zones via incorporation of the increase in stellar luminosity, which exclusively accounts for abiotic factors (see, e.g., Kasting et al., 1993; Rushby et al., 2013; Gallet et al., 2017; Tuchow and Wright, 2023), and not biological/technological feedbacks.
The outline of the paper is as follows. We present the methods and accompanying results pertaining to the impacts of global heating (from energy production) on putative biospheres (and technospheres) and climates in Section 2, where it is shown that the loss of habitable conditions might occur in $\sim 10^{3}$ yr. Next, we elucidate their ramifications for the potential trajectories of technological species (e.g., collapse; slowdown or cessation of growth; expansion into space) and detecting technosignatures in Section 3.
@@ 2.1 Thermodynamics of energy production
The Second Law of thermodynamics dictates that whenever energy $E$ is converted into useful work $W$, waste heat $\mathcal{Q}$ is produced and dissipated in the environment (e.g., Summers, 1971; Struchtrup, 2024; Blundell and Blundell, 2010; Schroeder, 2020). If the conversion mechanism exhibits an efficiency of $\eta=W/E$, the amount of energy that goes into heating is expressible as
@@
Carnot’s theorem canonically sets a theoretical limit on the maximum efficiency attainable by a cyclical process operating between temperatures $T_{C}$ and $T_{H}$ (e.g., Struchtrup, 2024; Schroeder, 2020), as given by the famous formula:
@@
It is, therefore, unavoidable that energy conversion (to thermodynamic work) enabling technological activity produces a finite amount of waste heat, $\mathcal{Q}_{\rm tech}$ (Budyko, 1969, 1972). When energy is extracted from sources that would otherwise be rendered “inactive” (such as burning fossil fuels or exploiting nuclear reactions), $\mathcal{Q}_{\rm tech}$ will constitute an extra heat source that is fed into the planetary system, in addition to the “natural” heat deposited by stellar irradiation or by the dissipation of “intrinsic” (e.g., radioactivity) planetary energy sources. It is worth recognizing that the warming connected to $\mathcal{Q}_{\rm tech}$ is conceptually different from that caused by the enhanced (viz., anthropogenic) greenhouse effect resulting from alterations of atmospheric chemistry. While in principle a “zero emissions” energy conversion strategy ought to be viable, the production of waste heat can only be minimized, but never completely eliminated.
@@
As underscored in the preceding exposition, the waste heat is comparable to the input energy for known sources. Therefore, we will hereafter adopt the ostensibly reasonable premise that the waste heat generated is roughly equal to the energy consumption associated with technological activity, unless indicated otherwise. If we postulate that the magnitude of waste heat generated is diminished by about an order of magnitude, this does not alter our qualitative findings, and it modifies the quantitative results by merely a modest degree; this facet is implicit in the paragraph below (17).
@@ 2.2 Global heating from technological activity
At equilibrium, the incoming stellar energy reaching a planetary surface has to equal the outgoing flux emitted by the planet itself. The basic (zero-dimensional and equilibrium) energy balance equation can be expressed as (e.g., Jacob 1999, Chapter 7, Vallis 2012, Chapter 1, and Tziperman 2022, Chapter 2):
@@
where $S$ is the stellar flux incident at the top of the atmosphere, $A$ is the planetary albedo, $\varepsilon$ is the fraction of terrestrial infrared radiation absorbed by the atmosphere (a parameter that can be calibrated to model greenhouse warming), $T$ is the planetary surface temperature (which is not the same as the planet’s equilibrium blackbody temperature), and $Q_{\rm tech}$ is the waste heat flux (units of power per unit area) ensuing from technological activity. Where needed (i.e., not for every calculation), we will henceforth adopt an average albedo of $A\approx 0.3$ for modern Earth (Trenberth et al., 2009; Stephens et al., 2015), as well as for “Earth-like” habitable worlds (i.e., Earth-analogs) (cf. Del Genio et al., 2019; Rushby et al., 2019; Madden and Kaltenegger, 2020). [^4] However, as long as $A$ is not very close to unity (i.e., $A\rightarrow 1$), our forthcoming quantitative results are not particularly sensitive to this parameter.
It should, however, be noted that (3) neglects the supplemental role of convection (Hartmann, 2016; Tziperman, 2022), which may be of importance for certain temperate planets orbiting M-dwarfs (Shields et al., 2016). Likewise, we neglect internal heat sources in (3), which would enter as additional terms on the LHS, and would serve to elevate the temperature further, and might possibly diminish the timescales computed later in this paper. The possible sources are radioactivity and tidal heating, both of which are relatively modest on Earth compared to $S$; the former translates to a heat flux of $\sim 7.4\times 10^{-2}$ Wm-2 (The Borexino Collaboration, 2020). On some habitable worlds, however, it is plausible that tidal heating (Jackson et al., 2008; Barnes et al., 2009; Driscoll and Barnes, 2015; Dobos et al., 2019) and radiogenic heating (Frank et al., 2014; Lingam and Loeb, 2020; Nimmo et al., 2020) could be higher, and thus necessitate inclusion in the global energy budget equation.
The energy balance in (3) is applied at the planetary surface. The first and the second terms on the LHS account for incoming stellar radiation and the thermal radiation emitted by the atmosphere (toward the surface), whereas the RHS is the outgoing thermal radiation from the planetary surface. The last term on the LHS is the extra contribution arising from technological activity. For the sake of the subsequent exposition, we rewrite (3) in terms of the parameter $\tau\equiv 1-\varepsilon/2$, which has been interpreted as an effective “transmittance” in some publications (e.g., Steiglechner, 2018), thereupon simplifying to yield
@@
where it should be noted that $\tau$ is innately dependent on atmospheric properties, such as its composition, and encapsulates a fair amount of complexity.
All values should be inherently perceived as global averages. As mentioned in Section 2.1, $Q_{\rm tech}$ is anticipated to be comparable to the total power converted to support technological activity. As a reference, adopting a global primary energy consumption of $1.8\times 10^{5}$ TWh in 2022 (Ritchie et al., 2024), [^5] https://ourworldindata.org/energy-production-consumption and averaging over the entire surface of the Earth ($5.1\times 10^{14}\ \rm m^{2}$) yields an estimate of $Q_{\rm tech}\sim 0.04$ Wm-2, which matches the calculation reported in Kleidon et al. (2023), acting as a consistency check. This is currently a negligible fraction of the radiative forcing attributable to the anthropogenic greenhouse effect (computed to be $2.7$ Wm-2) responsible for global warming (Kleidon et al., 2023).
@@
where we have substituted (5) into (6). An interesting feature of the above equation is that the lower bound on $\Delta T/T_{0}$ is attained in the low-albedo limit of $A\rightarrow 0$, and is given by
@@
It must be recalled that our goal in this work is not to perform state-of-the-art climate modeling, but to formulate a heuristic treatment that captures the salient trends (in consideration of the many variables at play). Furthermore, as intimated in Section 2.1 and particularly Section 2.4, the supposition of exponential growth in power consumption and waste heat generation ensures that most of our predictions pertaining to the timescales (of various variables) are quite robust, owing to a weak (i.e., logarithmic) dependence.
@@ 2.3 Global heating from stellar energy conversion
Hitherto, $Q_{\rm tech}$ has been interpreted as extra heating, in addition to that naturally produced by stellar radiation, from a source that would otherwise be sequestered within the planet (e.g. by fossil fuel burning or nuclear reactions). We will now analyze the case wherein energy for technological activity is entirely produced via suitable (e.g., photovoltaic) conversion of the stellar flux reaching the planetary surface (refer to De Vos 2008; Weinstein et al. 2015; Khan and Arsalan 2016; Kabir et al. 2018; Gong et al. 2019; Hayat et al. 2019).
Ultimately, an appropriate fraction of such energy converted into useful work (for instance, through the production of electricity) is dissipated as heat. At first, it may appear that this energy usage ought not translate to an elevation of the planetary surface temperature, since all the incoming and outgoing energy is already seemingly included in (4) with $Q_{\rm tech}=0$. However, the efficient conversion of stellar energy necessitates the collection of incident flux that would otherwise be reflected to space by tapping into methods such as surface texturing (Narasimhan and Cui, 2013; Kim et al., 2020b) and anti-reflective coatings (Raut et al., 2011; Sarkın et al., 2020; Shanmugam et al., 2020), which thereby can induce a net decrease of the planetary albedo (Berg et al., 2015; Barron-Gafford et al., 2016; Broadbent et al., 2019; Ziar et al., 2019; Sailor et al., 2021; Xu et al., 2024). [^7] We caution, however, that solar energy conversion by means of photovoltaic cells engenders an intricate cascade of climatic effects on Earth currently, which may contribute to either regional heating (Barron-Gafford et al., 2016; Broadbent et al., 2019; Sailor et al., 2021; Sun et al., 2022; Khan and Santamouris, 2023) or cooling (Masson et al., 2014; Hu et al., 2016; Zhang and Xu, 2020; Xu et al., 2024); by the same token, future extrapolations to extensive deployment of these cells could have major attendant uncertainties. If $f_{s}$ is the fractional area covered by solar panels and $A_{s}$ is the panel albedo, the balance equation becomes:
@@
The above result may be viewed as a lower bound because it only comprises the heating from change(s) in albedo, and does not incorporate the finite efficiency of photovoltaics (Kirchartz and Rau, 2018; Kim et al., 2020a; Green et al., 2023), [^8] For a single junction photovoltaic cell, Shockley and Queisser (1961) derived a theoretical efficiency limit, which has been subsequently modified both empirically and theoretically (Nelson, 2003; Krogstrup et al., 2013; Rühle, 2016; Ehrler et al., 2020; Markvart, 2022). and further losses during energy conversion. The fraction of stellar flux converted into usable energy flux $\mathcal{P}$ (e.g., by way of photovoltaics) is
@@
In the idealized limit of $A_{s}=0$, the contribution to heating can decrease by a factor of $A$ for the same energy consumption (i.e., a fixed value of $\mathcal{P}$), as revealed from inspecting (14); this case is consistent with the simplified results presented in Berg et al. (2015, Section 4). The maximum power attainable via photovoltaic cells (assuming an unrealistic 100% efficiency in stellar energy conversion) corresponds to the case with $f_{s}=1$ and $A_{s}=0$ in (13), to wit, when all the stellar flux reaching the entire planetary surface ($S/4$) is collected by panels, without any reflection to space or miscellaneous losses. In this situation, the planet is anticipated to experience heating by an extra $A\times(S/4)$, as revealed by (14). On plugging these values in (11), it can be shown that $T$ for this planet corresponds to that of a perfect absorber, as expected.

### v2

@@ Abstract
Waste heat production represents an inevitable consequence of energy conversion as per the laws of thermodynamics. Based on this fact, by using simple theoretical models, we analyze constraints on the habitability of Earth-like terrestrial planets hosting putative technological species and technospheres characterized by persistent exponential growth of energy consumption and waste heat generation. In particular, we quantify the deleterious effects of rising surface temperature on biospheric processes and the eventual loss of liquid water. Irrespective of whether these sources of energy are ultimately stellar or planetary (e.g., nuclear, fossil fuels) in nature, we demonstrate that the loss of habitable conditions on such terrestrial planets may be expected to occur on timescales of $\lesssim 1000$ years, as measured from the start of the exponential phase, provided that the annual growth rate of energy consumption is of order $1\%$. We conclude with a discussion of the types of evolutionary trajectories that might be feasible for industrialized technological species, and we sketch the ensuing implications for technosignature searches.
@@ 1 Introduction
It is now firmly established that the upsurge in atmospheric CO2 concentration caused by human technological activities has been elevating Earth’s average surface temperature (e.g., Archer, 2016; Hartmann, 2016; Dessler, 2021; Magnan et al., 2021; Abbass et al., 2022; Romm, 2022; Tziperman, 2022). Nevertheless, it is worth appreciating that ongoing anthropogenic warming is not exclusively a consequence of the intensified greenhouse effect stemming from fossil fuel utilization (though the latter phenomenon is unequivocally proven to be the chief driver). Instead, as elucidated in a number of analyses and overviews (e.g., Chaisson, 2008; Flanner, 2009; Kleidon, 2012; Ming et al., 2014; Lu et al., 2017; Kleidon et al., 2023), this facet of anthropogenic warming embodies a broader outcome of energy production and conversion that underpins modern (viz., globalized and industrialized) human societies. The pioneering 1969 publication by the Soviet scientist Mikhail I. Budyko explicitly underscored this theme (Budyko, 1969, pg. 618):
@@
All the energy used by man is transformed into heat, the main portion of this energy being an additional source of heat as compared to the present radiation gain. Simple calculations show (Budyko, 1961) that with the present rate of growth of using energy the heat produced by man in less than two hundred years will be comparable with the energy coming from the sun.
@@
This feature arises from the fundamental fact that energy transduction inevitably generates waste heat (e.g., Budyko, 1969; Summers, 1971; Smil, 2008) [^1] Interestingly, around the same time as Budyko’s paper, Larry Niven’s classic science fiction novel Ringworld (1970) also acknowledged the importance of waste heat as an evolutionary factor for technological species (Niven, 1970).. Although the current contribution of this factor to global warming on Earth is negligible compared to the impact of carbon emissions (Ming et al., 2014; Kleidon et al., 2023), [^2] In localized regions endowed with high industrialization (e.g., large and dense cities), however, it is already a notable contributor to warming (Ichinose et al., 1999; Block et al., 2004; Ohashi et al., 2007; Allen et al., 2011; Zhang et al., 2013; Bohnenstengel et al., 2014; Salamanca et al., 2014; Stewart and Kennedy, 2017; Sun et al., 2018; Raj et al., 2020; Molnár et al., 2020), and this effect is anticipated to increase in the future. it is expected to amplify quite substantially in the next century in the event that the current growth rate of global energy consumption persists (Lu et al., 2017). The consequent rise in global temperature (just from this process) derived from assorted climate models may reach up to $\sim 1$ K in this duration (Flanner, 2009; Nickolaenko, 2009; Steiglechner et al., 2019).
@@
Cruder estimates, when extrapolated over longer intervals, project a temperature increase of $\sim 10$ K over a period of $350$ years, with the assumption of energy trends similar to those of today (Chaisson, 2008). Likewise, a cognate analysis by Mullan and Haqq-Misra (2019) (see also von Hoerner 1975; Dutil and Dumas 2009) concluded that direct heating engendered by progressively higher energy consumption consistent with current rates could dramatically impede human habitability on timescales of merely hundreds of years in the future. It is crucial to realize that this contribution to global warming would persist even in the scenario of achieving a complete transition to zero CO2-emissions energy production.
@@
While this effect is of profound importance for the long-term future and fate of human civilization, it also has broader significance in the general context of planetary evolution and the potential for harboring technologically capable life in the universe – this latter domain is closely linked with the detection of technosignatures, a burgeoning area of astrobiology (Ćirković, 2012; Lingam and Loeb, 2021; Socas-Navarro et al., 2021), with roots traceable to the 1960s (Sagan, 1963; Drake, 1965; Shklovskii and Sagan, 1966; Tarter, 2001). To be specific, we are free to inquire whether energetic growth, often innately presumed to be associated with technological activity (Kardashev, 1985), imposes a universal fundamental limit on the habitability of a planetary environment, arising simply from thermodynamic considerations, as indicated in von Hoerner (1975).
@@
Thus, the summum bonum of this paper is to perform a preliminary investigation of how the habitability of Earth-like planets would be affected by the presence of life specifically capable of engaging in progressively energy-demanding technological activities. One of our primary objectives is to gauge the maximum lifespan of habitable planets that harbor an evolving “technosphere” (Donges et al., 2017; Zalasiewicz et al., 2017; Herrmann-Pillath, 2018), by making certain assumptions regarding the growth of energy consumption by an industrialized technological species. This analysis parallels the canonical approach used to estimate the lifetime of continuously habitable zones via incorporation of the increase in stellar luminosity, which exclusively accounts for abiotic factors (see, e.g., Kasting et al., 1993; Rushby et al., 2013; Gallet et al., 2017; Tuchow and Wright, 2023), and not biological/technological feedbacks.
The outline of the current paper is as follows. We present the methods and accompanying results that pertain to the impacts of global heating (from energy production) on putative biospheres (and technospheres) and climates in Section 2, where it is shown that the loss of habitable conditions might occur in $\sim 10^{3}$ yr. Next, we elucidate their ramifications for the potential trajectories of technological species (e.g., collapse; slowdown or cessation of growth; expansion into space) and the detection of technosignatures in Section 3.
@@ 2.1 Thermodynamics of energy production
The Second Law of thermodynamics dictates that whenever energy $E$ is converted into useful work $W$, waste heat $\mathcal{Q}$ is produced and dissipated in the environment (e.g., Summers, 1971; Reichl, 1998; Blundell and Blundell, 2010; Struchtrup, 2024). If the conversion mechanism exhibits an efficiency of $\eta=W/E$, the amount of energy that goes into heating is expressible as
@@
Carnot’s theorem canonically sets a theoretical limit on the maximum efficiency attainable by a cyclical process operating between temperatures $T_{C}$ and $T_{H}$ (e.g., Schroeder, 2020; Struchtrup, 2024), as given by the famous formula:
@@
It is, therefore, unavoidable that energy conversion (to thermodynamic work) that enables technological activity produces a finite amount of waste heat, $\mathcal{Q}_{\rm tech}$ (Budyko, 1969, 1972). When energy is extracted from sources that would otherwise be rendered “inactive” (such as burning fossil fuels or exploiting nuclear reactions), $\mathcal{Q}_{\rm tech}$ will constitute an extra heat source that is fed into the planetary system, in addition to the “natural” heat deposited by stellar irradiation or by the dissipation of “intrinsic” planetary energy sources (e.g., radioactivity). It is worth recognizing that the warming connected to $\mathcal{Q}_{\rm tech}$ is conceptually different from that caused by the enhanced (viz., anthropogenic) greenhouse effect that results from alterations of atmospheric chemistry. While in principle a “zero emissions” energy conversion strategy ought to be viable, the production of waste heat can only be minimized, but never completely eliminated.
@@
As underscored in the preceding exposition, the waste heat is comparable to the input energy for known sources. Therefore, we will hereafter adopt the ostensibly reasonable premise that the waste heat generated is roughly equal to the energy consumption associated with technological activity, unless indicated otherwise. If we postulate that the magnitude of waste heat generated is diminished by about an order of magnitude, this does not alter our qualitative findings, and it modifies the quantitative results by merely a modest degree; this is implicit in the paragraph below (17).
@@ 2.2 Global heating from technological activity
At equilibrium, the incoming stellar energy that reaches a planetary surface has to equal the outgoing flux emitted by the planet itself. The basic (zero-dimensional and equilibrium) energy balance equation can be expressed as (e.g., Jacob 1999, Chapter 7, Vallis 2012, Chapter 1, and Tziperman 2022, Chapter 2):
@@
where $S$ is the stellar flux incident at the top of the atmosphere, $A$ is the planetary albedo, $\varepsilon$ is the fraction of terrestrial infrared radiation absorbed by the atmosphere (a parameter that can be calibrated to model greenhouse warming), $T$ is the planetary surface temperature (which is not the same as the planet’s equilibrium blackbody temperature), and $Q_{\rm tech}$ is the waste heat flux (units of power per unit area) that ensues from technological activity. Where needed (i.e., not for every calculation), we will henceforth adopt an average albedo of $A\approx 0.3$ for modern Earth (Trenberth et al., 2009; Stephens et al., 2015), as well as for “Earth-like” habitable worlds (i.e., Earth-analogs) (cf. Del Genio et al., 2019; Rushby et al., 2019; Madden and Kaltenegger, 2020). [^4] However, as long as $A$ is not very close to unity (i.e., $A\rightarrow 1$), our forthcoming quantitative results are not particularly sensitive to this parameter.
It should, however, be noted that (3) neglects the supplemental role of convection (Hartmann, 2016; Tziperman, 2022), which may be of importance for certain temperate planets orbiting M-dwarfs (Shields et al., 2016). Likewise, we neglect internal heat sources in (3), which would enter as additional terms on the left-hand side (LHS), and would serve to elevate the temperature further, and might possibly diminish the timescales computed in this paper. The possible sources are radioactivity and tidal heating, both of which are relatively modest on Earth compared to $S$; the former translates to a heat flux of $\sim 7.4\times 10^{-2}$ Wm-2 (The Borexino Collaboration, 2020). On some habitable worlds, however, it is plausible that tidal heating (Jackson et al., 2008; Barnes et al., 2009; Driscoll and Barnes, 2015; Dobos et al., 2019) and radiogenic heating (Frank et al., 2014; Lingam and Loeb, 2020; Nimmo et al., 2020) could be higher and thus necessitate inclusion in the global energy budget equation.
The energy balance in (3) is applied at the planetary surface. The first and the second terms on the LHS account for incoming stellar radiation and the thermal radiation emitted by the atmosphere (toward the surface), whereas the right-hand side (RHS) is the outgoing thermal radiation from the planetary surface. The last term on the LHS is the extra contribution arising from technological activity. For the sake of the subsequent exposition, we rewrite (3) in terms of the parameter $\tau\equiv 1-\varepsilon/2$, which has been interpreted as an effective “transmittance” in some publications (e.g., Steiglechner, 2018), thereupon simplifying to yield
@@
where it should be noted that $\tau$ is innately dependent on atmospheric properties, such as atmospheric composition, and encapsulates a fair amount of complexity.
All values should be inherently perceived as global averages. As mentioned in Section 2.1, $Q_{\rm tech}$ is anticipated to be comparable to the total power converted to support technological activity. As a reference, adopting a global primary energy consumption of $1.8\times 10^{5}$ TWh in 2022 (Ritchie et al., 2024), [^5] https://ourworldindata.org/energy-production-consumption and averaging over the entire surface of Earth ($5.1\times 10^{14}\ \rm m^{2}$) yields an estimate of $Q_{\rm tech}\sim 0.04$ Wm-2, which matches the calculation reported in Kleidon et al. (2023), acting as a consistency check. This is currently a negligible fraction of the radiative forcing attributable to the anthropogenic greenhouse effect (computed to be $2.7$ Wm-2) responsible for global warming (Kleidon et al., 2023).
@@
where we have substituted (5) into (6). An interesting feature of the above equation is that the lower bound on $\Delta T/T_{0}$ is attained in the low-albedo limit of $A\rightarrow 0$ and is given by
@@
It must be recalled that our goal in this work is not to perform state-of-the-art climate modeling but to formulate a heuristic treatment that captures the salient trends (in consideration of the many variables at play). Furthermore, as intimated in Section 2.1 and particularly Section 2.4, the supposition of exponential growth in power consumption and waste heat generation ensures that most of our predictions pertaining to the timescales (of various variables) are quite robust, owing to a weak (i.e., logarithmic) dependence.
@@ 2.3 Global heating from stellar energy conversion
Hitherto, $Q_{\rm tech}$ has been interpreted as extra heating, in addition to that naturally produced by stellar radiation, from a source that would otherwise be sequestered within the planet (e.g. by fossil fuel burning or nuclear reactions). We will now analyze the case wherein energy for technological activity is entirely produced via suitable (e.g., photovoltaic) conversion of the stellar flux that reaches the planetary surface (refer to De Vos 2008; Weinstein et al. 2015; Khan and Arsalan 2016; Kabir et al. 2018; Gong et al. 2019; Hayat et al. 2019).
Ultimately, an appropriate fraction of such energy converted into useful work (for instance, through the production of electricity) is dissipated as heat. At first, it may appear that this energy usage ought not to translate to an elevation of the planetary surface temperature, since all the incoming and outgoing energy is already seemingly included in (4) with $Q_{\rm tech}=0$. However, the efficient conversion of stellar energy necessitates the collection of incident flux that would otherwise be reflected to space by tapping into methods such as surface texturing (Narasimhan and Cui, 2013; Kim et al., 2020b) and anti-reflective coatings (Raut et al., 2011; Sarkın et al., 2020; Shanmugam et al., 2020), which thereby can induce a net decrease of the planetary albedo (Berg et al., 2015; Barron-Gafford et al., 2016; Broadbent et al., 2019; Ziar et al., 2019; Sailor et al., 2021; Xu et al., 2024). [^7] We caution, however, that solar energy conversion by means of photovoltaic cells engenders an intricate cascade of climatic effects on Earth currently, which may contribute to either regional heating (Barron-Gafford et al., 2016; Broadbent et al., 2019; Sailor et al., 2021; Sun et al., 2022; Khan and Santamouris, 2023) or cooling (Masson et al., 2014; Hu et al., 2016; Zhang and Xu, 2020; Xu et al., 2024); by the same token, future extrapolations to extensive deployment of these cells could have major attendant uncertainties. If $f_{s}$ is the fractional area covered by solar panels and $A_{s}$ is the panel albedo, the balance equation becomes:
@@
The above result may be viewed as a lower bound because it only comprises the heating from change(s) in albedo and does not incorporate the finite efficiency of photovoltaics (Kirchartz and Rau, 2018; Kim et al., 2020a; Green et al., 2023) [^8] For a single junction photovoltaic cell, Shockley and Queisser (1961) derived a theoretical efficiency limit, which has been subsequently modified both empirically and theoretically (Nelson, 2003; Krogstrup et al., 2013; Rühle, 2016; Ehrler et al., 2020; Markvart, 2022). and further losses during energy conversion. The fraction of stellar flux converted into usable energy flux $\mathcal{P}$ (e.g., by way of photovoltaics) is
@@
In the idealized limit of $A_{s}=0$, the contribution to heating can decrease by a factor of $A$ for the same energy consumption (i.e., a fixed value of $\mathcal{P}$), as revealed from inspecting (14); this case is consistent with the simplified results presented in Berg et al. (2015, Section 4). The maximum power attainable via photovoltaic cells (assuming an unrealistic 100% efficiency in stellar energy conversion) corresponds to the case with $f_{s}=1$ and $A_{s}=0$ in (13), to wit, when all the stellar flux that reaches the entire planetary surface ($S/4$) is collected by panels, without any reflection to space or miscellaneous losses. In this situation, the planet is anticipated to experience heating by an extra $A\times(S/4)$, as revealed by (14). On plugging these values in (11), it can be shown that $T$ for this planet corresponds to that of a perfect absorber, as expected.